OpenStudy (anonymous):

When using math induction to prove that (n^3 - n + 3) is divisible by 3 for all natural numbers. What equation would you need to show is divisible by 3?

OpenStudy (anonymous):

you can do this easier without induction by factoring

OpenStudy (anonymous):

since \(x^3-n+3=(n-1)\times n \times (n+1)+3\) since one of the three consecutive integers \(n-1, n, n+1\) is divisible by 3 and 3 is also

OpenStudy (anonymous):

or you can proceed by induction if you have to show it is true if \(n=1\) and then assume it is true for \(k\) i.e. assume \[k^3-k+3\] is divisible by 3, then prove if that is the case the \[(k+1)^3-(k+1)+3\] is also divisible by 3 it is a bunch of algebra to find the hidden \(k^3-k+3\) in \((k+1)^3-(k+1)+3\)

OpenStudy (anonymous):

Thank you!